OpenStudy (anonymous):

1. Translate the words into a mathematical expression: n more than 8. 2. Translate the words into a mathematical expression: the product of 4 and a number g. 3. Translate the words into a mathematical expression: the difference of 5 and twice x. 4. Translate the words into a mathematical expression: the quotient of 6 and x. (Points : 1) 5. Translate the words into a mathematical expression: the sum of 3 times m and 9 6.Translate the words into a mathematical expression: 7 plus the quantity of 8 minus 6. 7. Evaluate 8 - n if n = 3 8. Simplify (20 - 2) ÷ 9. 9.Evaluate 3x + 1 if x = 4.

OpenStudy (anonymous):

is this just your homework

OpenStudy (anonymous):

yes its my homework i need it to raise my grade to go to public school please help meeeee

OpenStudy (anonymous):

1. 8+n 2. 4g 3. 2x-5 4. 6/x 5. 3m +9 6. 7+(8-6) 7. plug in. you get this. 8-3 which equals 5. Your answer is: 5 8. (20 - 2) ÷ 9 = 18 ÷ 9 = 2 Your answer is: 2 9. 3x + 1 if x = 4. again, plug in the value of x wherever an x is present. so you get this. 3(4)+1=12+1=13 Your answer is 13.

OpenStudy (anonymous):

remember you need to read the terms and the way they group and write everything carefully. :) that's how you do the ones where you translate into mathematical terms :) and as for the others, you plug in :) do they make sense?? :)

OpenStudy (anonymous):

yes thank you! :)

OpenStudy (anonymous):

ur welkers :) glad to help!

OpenStudy (anonymous):

have fun in public school! lol :)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

2. Five is a rational number. (Points : 1) True False 3. Mark every set that applies in the remaining questions. Classify -5. (Points : 1) Whole Number Integer Rational Number Irrational Number 4. Classify (Points : 1) Whole Number Integer Rational Number Irrational Number 5. Classify (Points : 1) Whole Number Integer Rational Number Irrational Number 6. Classify 8. (Points : 1) Whole Number Integer Rational Number Irrational Number 7. Classify -3.4 (Points : 1) Whole Number Integer Rational Number Irrational Number 8. Classify (Points : 1) Whole Number Integer Rational Number Irrational Number 9. Classify 0.626539212... (Points : 1) Whole Number Integer Rational Number Irrational Number 10. Classify 0.767676... (Points : 1) Whole Number Integer Rational Number Irrational Number

OpenStudy (anonymous):

number 8 is classify 36 squared

OpenStudy (anonymous):

and number 5 is classify 5 squared

OpenStudy (anonymous):

and number 4 classify negative one fifth

OpenStudy (anonymous):

2. Should be TRUE (b/c it's rational as a fraction 5/1 so it should be as 5 as well) 3. not 100% sure, but should be whole number, integer and rational number 4. irrational number 5. whole number, integer, rational number 6. whole number, integer, rational number 7. rational number 8. whole number, integer, rational number 9. irrational number 10. irrational number

OpenStudy (anonymous):

okay, so idk if they are completely 100% correct... but most should be correct.. sorry if some are wrong tho!! :O i really hope most are correct tho!! they should be :) if u read the definitions of those terms, they should match up :)

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

np! :)